# Learning to Group and Grasp Multiple Objects

### v1

@@ IV Network Selection and Learning
We use the method proposed in [3] for imitation learning. The method enables the prediction of actions not just for a single step but for multiple future steps based on past observations. It can handle scenarios where multiple strategies exist and can achieve a goal with flexibility. Using the notations defined in the previous section, the input, prediction output, and usage of the model are as follows.
@@
Based on the past $T_{o}$ time steps of observations at time $t$ $\boldsymbol{O}_{t}=\{\boldsymbol{o}_{t-T_{o}+1}$, $\boldsymbol{o}_{t-T_{o}+2}$, ..., $\boldsymbol{o}_{t}\}$, the used method will infer actions for $T_{p}$ time steps, namely $\boldsymbol{A}_{t}=\{\boldsymbol{a}_{t-T_{o}+1}$, $\boldsymbol{a}_{t-T_{o}+2}$, ..., $\boldsymbol{a}_{t}$, $\boldsymbol{a}_{t+1}$, ..., $\boldsymbol{a}_{t+T_{p}-T_{o}}\}$. The inferred $\boldsymbol{A}_{t}$ includes the predicted past actions $\{\boldsymbol{a}_{t-T_{o}+1}$, $\boldsymbol{a}_{t-T_{o}+2}$, ..., $\boldsymbol{a}_{t-1}\}$ and the predicted future actions $\{\boldsymbol{a}_{t}$, $\boldsymbol{a}_{t+1}$, ..., $\boldsymbol{a}_{t+T_{p}-T_{o}}\}$. The predicted past actions correspond to the observations of the past $T_{o}$ time steps. The predicted future actions are the learned policies for the next $T_{p}-T_{o}$ time steps. The robot will execute $T_{a}$ steps of actions $\{\boldsymbol{a}_{t},\boldsymbol{a}_{t+1},...,\boldsymbol{a}_{t+T_{a}-1}\}$ for collecting new observations and conduct new predictions. $T_{a}$ is selected to be a value smaller than $T_{p}-T_{o}$. Fig. 3 illustrates the details of the time step division for observation, prediction, and execution.
Figure 3: The imitation learning method employed in this study takes as input the observations from the past $T_{o}$ time steps. The output consists of the inferred likely actions for the past $T_{o}$ time steps and predicted actions for the future $T_{p}-T_{o}$ time steps. The first $T_{a}$ predicted future actions are executed for iterative observation-prediction looping.
@@
The internal structure of the used method is illustrated in Fig. 4. There are in total $T_{o}$ observations. Each observation consists of two image frames respectively captured from the global and local cameras. They are processed separately using two ResNet-18 for feature extraction. Each image frame is compressed into a $N$-dimensional feature vector after feature extraction, and the total observations form two $N\times T_{o}$-dimensional vectors. Additionally, the observation includes the robot’s state, which is an $8$-dimensional vector for each observation as discussed in the previous section. In total, the robot states form a $8\times T_{o}$-dimensional vector. The two $N\times T_{o}$-dimensional vectors together with the $8\times T_{o}$-dimensional robot state vector are concatenated to form a $(2N+8)\times T_{o}$-dimensional vector which serves as the observation embedding of a U-Net core used in the diffusion policy network.
@@
The U-Net core takes a noisy action sequence (represented as a $8\times T_{p}$-dimensional vector) as input and predicts the noise in the representation. The output of the U-Net is another $8\times T_{p}$-dimensional vector where each value corresponds to the estimated noise in the input sequence. This predicted noise is subsequently subtracted from the input sequence within the denoising box shown in Fig. 4. The denoising substraction is mathematically
$\boldsymbol{A}_{t}^{k-1}=\alpha(\boldsymbol{A}_{t}^{k}-\gamma{\boldsymbol{\varepsilon_{\theta}}}(\boldsymbol{A}_{t}^{k},\boldsymbol{O}_{t},k)+\mathcal{N}(0,\sigma^{2}\boldsymbol{I})),$ (7)
@@
where $k$ indicates the $k$th denoising step. $\boldsymbol{A}_{t}^{k}$ represents the noisy action sequence at the $k$th denoising step. $\boldsymbol{\varepsilon_{\theta}}()$ represents the noise prediction network (namely the U-Net core). $\theta$ indicates the parameters of the network. For each denoising iteration, the iteration step index $k$ is encoded using a $P$-dimensional sinusoidal positional embedding and is also incorporated into the U-Net core to provide temporal context. The positional embedding helps guide the network in estimating how much noise remains to be removed. $\alpha$, $\gamma$, $\sigma$ are parameters for controlling the learning stability Note 1 Note 1 Note 1The default values from the LeRobot source code was adopted. Check the following webpage for details. https://github.com/huggingface/lerobot. The noisy action sequence is initialized by sampling a Gaussian noise distribution. After undergoing $K$ iterative denoising steps, the diffusion policy network will output the denoised $\boldsymbol{A}_{t}^{0}$ as predicted action sequences.
Figure 4: Internal mechanism of the diffusion policy network and its input-output processing. ResNet-18 is utilized to extract features from the input images. The policy network has a U-Net core, which accepts a noisy action sequence and predicts the noise within it. Image features and positional encoding are embedded into the U-Net core as conditional inputs. Through $K$ iterations, the network progressively removes the noise to acquire the action sequence most suitable for the observations.
@@
To train the denoising network, we draw a sequence of $\boldsymbol{O}_{t}$ and a sequence of expert action $\boldsymbol{A}_{t}^{0}$ from the collected data set, selects a random diffusion step $k$ and add Gaussian noises $\boldsymbol{\varepsilon}^{k}$ to the expert actions. The parameters $\theta$ of the U-Net core are learned by minimizing
@@ V Experiments and Analysis
V-A Experimental environment and parameter settings
[…]
For evaluation, we used two major categories of objects. Both of them have elongated rectangular shape. The first category consists of uniformly red 3D-printed objects, each with a length, width, and height of $120$ mm, $20$ mm, and $20$ mm. These objects are easily distinguishable, making them suitable for quickly verifying the effectiveness of the algorithm. The second category comprises real-world commercial products with similar shapes They include chocolate snack bars and fructose snack bars. The dimensions of these commercial products will be presented later in subsections. Based on the model trained using data collected from the first category, we further fine-tuned the model with the second category to evaluate its transferability and the effectiveness of the proposed method on real-world products.
@@ V-A Experimental environment and parameter settings
For our particular diffusion policy implementation, we set both the ResNet-18 output dimension $N$ and the positional encoding dimension $P$ to be 128, set $T_{o}$, $T_{a}$, and $T_{p}$ to be $2$, $8$, and $16$, respectively. That is, the learned policy takes the consecutive observations $\boldsymbol{o}_{t-1}$ and $\boldsymbol{o}_{t}$ as input and predicts future actions $\boldsymbol{a}_{t}$, $\boldsymbol{a}_{t+1}$, ..., $\boldsymbol{a}_{t+7}$ for the next 8 steps. By performing inference every 8 steps and executing the predicted actions in a closed-loop control manner, the system continuously adjusts its behavior. The hidden layers of the U-Net are set to $[512,1024,2024]$. The optimization method used is AdamW, with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. The batch size is set to $64$. For the diffusion process, the number of denoising steps is set to $K=100$.
V-B Experiment 1: Randomly placed 3D printed objects
In the first group of experiments, we focused on the influence by the number of randomly placed objects and number of training data. The evaluation is conducted on tasks where 3 to 6 randomly placed printed objects are grouped and transported using learned policies. The initial position of the UR3 end effector is set above the bin. The robot is expected to group the objects considering their distributions and then grasp and transport them to the bin.
During data collection, we conducted $200$ demonstrations, where the number of objects was randomly set between $3$ and $6$, and the target objects were randomly placed within the yellow frame. The collected training data were divided into two datasets of sizes $100$ and $200$, respectively, and a separate diffusion policy network was trained for each dataset to analyze the impact of data quantity on performance. Each diffusion policy network underwent $300000$ training steps.
During evaluation, $20$ trials were conducted for each number of placed objects. In each trial, the robot’s performance was assessed until it either successfully completed the task or failed. The following metrics were recorded for each trial.
  • $\triangleright$
    Number of grouping actions (#P): The number of grouping motions performed to gather the objects.
  • $\triangleright$
    Number of grasping actions (#G): The total number of grasp attempts.
  • $\triangleright$
    Sqeuence of grasped objects (Seq.)): The number of objects grasped by each grasping action, recorded in sequential order from the start of the task as a list.
  • $\triangleright$
    Task completion (C): Whether the all objects were successfully transported to the bin.
  • $\triangleright$
    Reason for task failure (R): If the task is not completed, the cause of failure is documented.
When the task fails to be completed, the failure causes can be attributed into the following three main reasons.
  • $\triangleright$
    Pushing failure (P): An object in contact during the pushing motion may deviate from the end effector and failed to be grouped.
  • $\triangleright$
    Grasping failure (G): The robot may fail to grasp the objects that were grouped together or drop the grasped object during transportation.
  • $\triangleright$
    Stagnation: The policy network may repeatedly suggest the same action and get into stagnation. If the robot remains stationary or repeated for more than 5 seconds, we consider it a failure. If stagnation occurs during grouping, it is classified as a pushing failure. Otherwise, it is classified as a grasping failure.
Table I shows the detailed results for all trials and statistics at the last row of each dataset-number of objects combination. Due to the limited space on the article page, this table contains numerous abbreviations. The specific meanings can be referenced in the notes at the bottom of the table. Especially in the statistical section, we computed the following values:
  • $\triangleright$
    Completion rate: The rate at which all objects were cleared over 20 trials.
  • $\triangleright$
    Transport rate: The average number of successfully transported objects divided by the number of objects placed across 20 trials.
  • $\triangleright$
    # Objects per grasp: The average number of objects transported per successful grasp across 20 trials.
The maximum values of these statistics across both trained networks and all numbers of objects are highlighted in lime, while the minimum values are highlighted in pink.
TABLE I: Results of randomly placed 3D printed objects
Data ID Number of Objects: 3 Number of Objects: 4 Number of Objects: 5 Number of Objects: 6
#P #G Seq. C R #P #G Seq. C R #P #G Seq. C R #P #G Seq. C R
100 1 2 2 (1,1) N P 2 2 (1,2) N P 1 1 (2) N P 0 0 (0) N P
2 2 2 (1,2) Y - 3 3 (1,1,2) Y - 4 4 (1,1,1,2) Y - 0 0 (0) N P
3 2 2 (2,1) Y - 0 0 (0) N P 0 0 (0) N P 0 0 (0) N P
4 0 0 (0) N P 0 0 (0) N P 3 3 (1,1,1) N P 0 0 (0) N P
5 2 2 (2,1) Y - 1 0 (0) N P 1 1 (1) N P 1 1 (1) N P
6 0 0 (0) N P 2 2 (2,2) Y - 1 1 (1) N P 0 0 (0) N P
7 2 2 (1,2) Y - 2 2 (2,2) Y - 0 0 (0) N P 2 2 (1,2) N P
8 0 0 (0) N P 1 1 (1) N P 0 0 (0) N P 2 2 (1,1) N P
9 1 1 (1) N P 2 2 (2,2) Y - 2 2 (3,2) Y - 1 1 (1) N P
10 2 2 (1,2) Y - 3 3 (2,1,1) Y - 2 2 (1,1) N P 1 1 (1) N P
11 1 1 (1) N P 0 0 (0) N P 1 1 (1) N P 0 0 (0) N P
12 1 1 (1) N P 3 3 (1,2,1) Y - 0 0 (0) N P 3 3 (3,2,1) Y -
13 1 1 (3) Y - 1 1 (1) N P 0 0 (0) N P 3 3 (1,3,2) Y -
14 1 1 (3) Y - 0 0 (0) N P 2 1 (3) N G 1 1 (3) N P
15 3 3 (1,1,1) Y - 3 3 (2,1,1) Y - 0 0 (0) N P 2 2 (1,2) N P
16 2 2 (2,1) Y - 3 3 (1,2,1) Y - 3 3 (1,3,1) Y - 4 4 (2,2,1,1) Y -
17 2 2 (1,2) Y - 2 2 (2,2) Y - 0 0 (0) N P 0 0 (0) N P
18 2 2 (2,1) Y - 0 0 (0) N P 0 0 (0) N P 3 3 (2,2,1) N P
19 1 1 (1) N P 4 4 (1,1,1,1) Y - 2 1 (1) N G 1 1 (1) N P
20 3 3 (1,1,1) Y - 1 1 (1) N P 0 0 (0) N P 5 5 (1,1,1,1,2) -
St Completion rate:      60% Completion rate:      50% Completion rate:      15% Completion rate:      10%
Transport rate:         70% Transport rate:         58% Transport rate:         29% Transport rate:         37%
# Objects per grasp:   1.4 # Objects per grasp:   1.44 # Objects per grasp:   1.45 # Objects per grasp:   1.52
200 1 2 2 (2,1) Y - 1 1 (2) N P 3 3 (1,3,1) Y - 3 3 (3,2,1) Y -
2 1 0 (0) N G 2 2 (3,1) Y - 4 4 (1,1,2,1) Y - 4 4 (1,2,2,1) Y -
3 2 2 (2,1) Y - 3 3 (2,1,1) Y - 3 3 (2,2,1) Y - 3 3 (2,3,1) Y -
4 1 1 (3) Y - 3 3 (2,1,1) Y - 3 3 (2,2,1) Y - 3 2 (1,3) N
5 2 2 (2,1) Y - 2 2 (2,2) Y - 2 2 (1,1) N P 3 3 (3,2,1) Y -
6 0 0 (0) N P 3 3 (1,2,1) Y - 4 4 (1,1,1,2) Y - 4 4 (1,2,1,2) Y -
7 3 3 (1,1,1) Y - 1 1 (3) N P 2 1 (1) N 5 5 (2,1,1,1,1) Y -
8 2 2 (1,2) Y - 4 4 (1,1,1,1) Y - 2 2 (1,1) N P 1 1 (3) N P
9 3 3 (1,1,1) Y - 2 2 (2,2) Y - 2 2 (3,2) Y - 3 3 (3,2,1) Y -
10 1 1 (3) Y - 3 3 (1,2,1) Y - 2 1 (1) N 4 4 (2,2,1,1) Y -
11 3 3 (1,1,1) Y - 0 0 (0) N P 4 4 (2,1,1,1) Y - 0 0 (0) N P
12 1 1 (3) Y - 3 3 (2,1,1) Y - 0 0 (0) N P 3 2 (2,1) N
13 1 0 (0) N G 2 2 (2,2) Y - 0 0 (0) N P 1 1 (2) N P
14 2 2 (1,1) N P 2 1 (2,) N G 3 3 (2,2,1) Y - 0 0 (0) N P
15 3 3 (1,1,1) Y - 3 3 (2,1,1) Y - 4 4 (1,1,2,1) Y - 4 4 (2,2,1,1) Y -
16 2 2 (2,1) Y - 2 2 (2,1) N P 1 1 (2) N P 4 4 (1,2,2,1) Y -
17 3 3 (1,1,1) Y - 4 4 (1,1,1,1) Y - 2 1 (2) N 1 1 (1) N P
18 2 2 (2,1) Y - 2 2 (1,3) Y - 4 4 (1,1,1,2) Y - 3 3 (2,1,3) Y -
19 1 1 (3) Y - 2 2 (1,1) N P 0 0 (0) N P 1 0 (0) N
20 2 2 (1,2) Y - 2 2 (3,1) Y - 4 4 (1,1,1,2) Y - 0 0 (0) N P
St Completion rate:      80% Completion rate:      70% Completion rate:      55% Completion rate:      55%
Transport rate:         83% Transport rate:         85% Trans. Rate: 65% Transport rate:         66%
# Objects per grasp:   1.43 # Objects per grasp:   1.51 # Objects per grasp:   1.42 # Objects per grasp:   1.68
  • ${}^{\text{Note}}$ Meanings of abbreviations: #P – Number of grouping actions; #G – Number of grasping actions; Seq. – Sequence of grasped objects by each grouping and grasping; C – Task completion (Y means Yes, all objects are cleared and transported. N means No, some objects remained due to failures); R – Reason for task failure (P means a pushing failure. G means a grasping failure); St – Statistical summary of the results above the row (Maximum values of the statistics are highlighted in lime, while minimum values are highlighted in pink).
We can see from the results that as the number of randomly placed graspable objects increases, the task completion rates decrease. Failures frequently occur during pushing and grouping. The frequency of the failures increases with the number of initially randomly placed objects. The majority of pushing failures are caused by stagnation. This may be attributed to the increasing variety in the input images as the number of objects grows, making it more difficult for the neural network to determine the optimal pushing strategy based on limited information. We can also see from the results that the sequence of grouping and grasping is quite random. The network chose different policies according to object distributions. Furthermore, the results indicate that an increase in training data has a positive impact on all metrics. Notably, when the number of objects reaches $6$, the robot grasps an average of 1.68 objects per attempt. This demonstrates that the policy network can effectively support multi-object grasping and thus help a robot improve picking efficiency.

### v3

@@ IV Network Selection and Learning
Based on the past $T_{o}$ time steps of observations at time $t$ $\boldsymbol{O}_{t}=\{\boldsymbol{o}_{t-T_{o}+1}$, $\boldsymbol{o}_{t-T_{o}+2}$, ..., $\boldsymbol{o}_{t}\}$, our diffusion policy is expect to infer actions for $T_{p}$ time steps, namely $\boldsymbol{A}_{t}=\{\boldsymbol{a}_{t-T_{o}+1}$, $\boldsymbol{a}_{t-T_{o}+2}$, ..., $\boldsymbol{a}_{t}$, $\boldsymbol{a}_{t+1}$, ..., $\boldsymbol{a}_{t+T_{p}-T_{o}}\}$. The inferred $\boldsymbol{A}_{t}$ includes the predicted past actions $\{\boldsymbol{a}_{t-T_{o}+1}$, $\boldsymbol{a}_{t-T_{o}+2}$, ..., $\boldsymbol{a}_{t-1}\}$ and the predicted future actions $\{\boldsymbol{a}_{t}$, $\boldsymbol{a}_{t+1}$, ..., $\boldsymbol{a}_{t+T_{p}-T_{o}}\}$. The predicted past actions correspond to the observations of the past $T_{o}$ time steps. The predicted future actions are the learned policies for the next $T_{p}-T_{o}$ time steps. The robot will execute $T_{a}$ steps of actions $\{\boldsymbol{a}_{t},\boldsymbol{a}_{t+1},...,\boldsymbol{a}_{t+T_{a}-1}\}$ for collecting new observations and conduct new predictions. $T_{a}$ is selected to be a value smaller than $T_{p}-T_{o}$. Fig. 3 illustrates the details of the time step division for observation, prediction, and execution.
@@
The internal structure of the used method is illustrated in Fig. 4. There are in total $T_{o}$ observations. Each observation consists of two image frames respectively captured from the global and local cameras. They are processed separately using two ResNet-18 for feature extraction. Each image frame is compressed into a $N$-dimensional feature vector after feature extraction, and the total observations form two $N\times T_{o}$-dimensional vectors. Additionally, the observation includes the robot’s state, which is an 888-dimensional vector for each observation as discussed in the previous section. In total, the robot states form a $8\times T_{o}$-dimensional vector. The two $N\times T_{o}$-dimensional vectors together with the $8\times T_{o}$-dimensional robot state vector are concatenated to form a $(2N+8)\times T_{o}$-dimensional vector which serves as the observation embedding of a U-Net core used in the diffusion policy network.
@@
The U-Net core takes a noisy action sequence (represented as a $8\times T_{p}$-dimensional vector) as input and predicts the noise in the representation. The output of the U-Net is another $8\times T_{p}$-dimensional vector where each value corresponds to the estimated noise in the input sequence. This predicted noise is subsequently subtracted from the input sequence within the denoising box shown in Fig. 4, which is mathematically
$\boldsymbol{A}_{t}^{k-1}=\alpha(\boldsymbol{A}_{t}^{k}-\gamma{\boldsymbol{\varepsilon_{\theta}}}(\boldsymbol{A}_{t}^{k},\boldsymbol{O}_{t},k)+\mathcal{N}(0,\sigma^{2}\boldsymbol{I})).$ (2)
@@
Here, $k$ indicates the $k$th denoising step. $\boldsymbol{A}_{t}^{k}$ represents the noisy action sequence at the $k$th denoising step. $\boldsymbol{\varepsilon_{\theta}}()$ represents the noise prediction network (namely the U-Net core). $\theta$ indicates the parameters of the network. For each denoising iteration, the iteration step index $k$ is encoded using a $P$-dimensional sinusoidal positional embedding and is also incorporated into the U-Net core to provide temporal context. The positional embedding helps guide the network in estimating how much noise remains to be removed. $\alpha$, $\gamma$, $\sigma$ are parameters for controlling the learning stability Note 2 Note 2 Note 2The default values from the LeRobot source code was adopted. Check the following webpage for details. https://github.com/huggingface/lerobot. The noisy action sequence is initialized by sampling a Gaussian noise distribution. After undergoing $K$ iterative denoising steps, the diffusion policy network will output the denoised $\boldsymbol{A}_{t}^{0}$ as predicted action sequences.
@@
To train the denoising network, we draw a sequence of $\boldsymbol{O}_{t}$ and a sequence of expert action $\boldsymbol{A}_{t}^{0}$ from the collected data set, select a random diffusion step $k$ and add Gaussian noises $\boldsymbol{\varepsilon}^{k}$ to the expert actions. The parameters $\theta$ of the U-Net core are learned by minimizing
@@
Figure 3: The diffusion policy takes as input the observations from the past $T_{o}$ time steps. It infers likely actions for the past $T_{o}$ time steps and predicts actions for the future $T_{p}-T_{o}$ time steps. The first $T_{a}$ predicted future actions are executed for iterative observation-prediction looping.
Figure 4: Internal mechanism of the diffusion policy network. It has a U-Net core that accepts a noisy action sequence and predicts the noise within it. Through $K$ iterations, the U-Net core progressively removes the noise to acquire the action sequence most suitable for the observations.
TABLE I: Results of randomly placed 3D-printed rectangular cuboids
ID Data = 100 Data = 200 Data = 300
# Objects: 4 # Objects: 6 # Objects: 4 # Objects: 6 # Objects: 4 # Objects: 6
#P #G Seq. C R #P #G Seq. C R #P #G Seq. C R #P #G Seq. C R #P #G Seq. C R #P #G Seq. C R
1 2 2 (1,2) N P 0 0 - N P 1 1 (2) N P 3 3 (3,2,1) Y - 2 2 (2,2) Y - 3 3 (3,1,2) Y -
2 3 3 (1,1,2) Y - 0 0 - N P 2 2 (3,1) Y - 4 4 (1,2,2,1) Y - 3 3 (1,2,1) Y - 4 4 (2,1,1,2) Y -
3 0 0 - N P 0 0 - N P 3 3 (2,1,1) Y - 3 3 (2,3,1) Y - 3 3 (1,2,1) Y - 4 4 (2,1,2,1) Y -
4 0 0 - N P 0 0 - N P 3 3 (2,1,1) Y - 3 2 (1,3) N G 2 2 (1,3) Y - 1 0 - N G
5 1 0 - N G 1 1 (1) N P 2 2 (2,2) Y - 3 3 (3,2,1) Y - 2 2 (3,1) Y - 1 1 (2) N P
6 2 2 (2,2) Y - 0 0 - N P 3 3 (1,2,1) Y - 4 4 (1,2,1,2) Y - 3 3 (1,2,1) Y - 4 4 (1,1,3,1) Y -
7 2 2 (2,2) Y - 2 2 (1,2) N P 1 1 (3) N P 5 5 (2,1,1,1,1) Y - 3 3 (2,1,1) Y - 3 3 (2,1,3) Y -
8 1 1 (1) N P 2 2 (1,1) N P 4 4 (1,1,1,1) Y - 1 1 (3) N P 0 0 - N P 3 3 (2,2,2) Y -
9 2 2 (2,2) Y - 1 1 (1) N P 2 2 (2,2) Y - 3 3 (3,2,1) Y - 2 1 (1,3) N G 3 3 (3,2,1) Y -
10 3 3 (2,1,1) Y - 1 1 (1) N P 3 3 (1,2,1) Y - 4 4 (2,2,1,1) Y - 3 3 (1,2,1) Y - 4 3 (2,2,1) N G
11 0 0 - N P 0 0 - N P 0 0 - N P 0 0 - N P 1 0 - N G 4 4 (1,3,1,1) Y -
12 3 3 (1,2,1) Y - 3 3 (3,2,1) Y - 3 3 (2,1,1) Y - 3 2 (2,1) N G 3 3 (1,2,1) Y - 1 1 (1) N P
13 1 1 (1) N P 3 3 (1,3,2) Y - 2 2 (2,2) Y - 1 1 (2) N P 3 3 (1,2,1) Y - 3 3 (2,2,2) Y -
14 0 0 - N P 1 1 (3) N P 2 1 (2) N G 0 0 - N P 2 2 (3,1) Y - 4 4 (1,1,3,1) Y -
15 3 3 (2,1,1) Y - 2 2 (1,2) N P 3 3 (2,1,1) Y - 4 4 (2,2,1,1) Y - 2 2 (2,2) Y - 3 3 (3,1,2) Y -
16 3 3 (1,2,1) Y - 4 4 (2,2,1,1) Y - 2 2 (2,1) N P 4 4 (1,2,2,1) Y - 3 3 (1,1,2) Y - 3 3 (2,2,2) Y -
17 2 2 (2,2) Y - 0 0 - N P 4 4 (1,1,1,1) Y - 1 1 (1) N P 3 3 (1,2,1) Y - 1 0 - N G
18 0 0 - N P 3 3 (2,2,1) N P 2 2 (1,3) Y - 3 3 (2,1,3) Y - 2 2 (1,3) Y - 3 3 (3,2,1) Y -
19 4 4 (1,1,1,1) Y - 1 1 (1) N P 2 2 (1,1) N P 1 0 - N G 3 3 (2,1,1) Y - 1 0 - N G
20 1 1 (1) N P 5 5 (1,1,1,1,2) Y - 2 2 (3,1) Y - 0 0 - N P 3 3 (2,1,1) Y - 3 3 (2,2,2) Y -
St 50%, 58%, 1.44 10%, 57%, 1.52 70%, 85%, 1.51 55%, 66%, 1.68 85%, 86%, 1.50 70%, 77%, 1.77
  • Note 1
    #P – Number of grouping actions; #G – Number of grasping actions; Seq. – Objects grasped per action; C – Task completion; St – Summary.
  • Note 2
    The three values in the St row are, in order: completion rate, delivery rate, and number of objects per grasp.
@@ V Experiments and Analysis
Our experiments were conducted using a Universal Robots UR3 manipulator with a Robotiq 2F-85 gripper. The target objects were placed within a yellow frame measuring 300 mm $\times$ 400 mm and are expected to be delivered to an adjacent tray. Two Intel RealSense D435 cameras are used as the global and local cameras. Only the RGB data was used and the image sizes were compressed to 240×320240$\times$ 320240 × 320 pixels to constrain GPU memory consumption. A PC with Intel Core Ultra 7 285K, DDR5-5600 192GB memory, and NVIDIA RTX A6000 GPU was used to train the diffusion policy. When collecting human expert data, another Intel RealSense D435 camera is used to track human hand poses and gripping states. The primary objects used in the experiments were 3D-printed rectangular cuboids with dimensions of 120 mm $\times$ 20 mm $\times$ 20 mm. To evaluate the generalization capability of the learned policy, we additionally prepared a set of shape variants with slightly altered aspect ratios as well as real-world consumer products, such as snack bars. In our particular diffusion policy implementation, we set both the ResNet-18 output dimension $N$ and the positional encoding dimension $P$ to be 128, set $T_{o}$, $T_{a}$, and $T_{p}$ to be 2, 8, and 16, respectively. That is, the learned policy takes the consecutive observations $\boldsymbol{o}_{t-1}$ and $\boldsymbol{o}_{t}$ as input and predicts future actions $\boldsymbol{a}_{t}$, $\boldsymbol{a}_{t+1}$, ..., $\boldsymbol{a}_{t+7}$ for the next 8 steps. By performing inference every 8 steps and executing the predicted actions in a closed-loop control manner, the system continuously adjusts its behavior. The hidden layers of the U-Net are set to [512, 1024, 2024]. The optimization method used is AdamW, with a learning rate of $10^{-4}$ and a weight decay of $10^{-6}$. The batch size is set to 64. For the diffusion process, the number of denoising steps is set to $K=100$.
V-A Experiment 1: Randomly placed 3D printed cuboids
First, we study the influence of training data scales. During data collection, we conducted 300300300 demonstrations, where the number of objects was randomly set between 333 and 666, and the target objects were randomly placed within the yellow frame. The collected training data were divided into three datasets of sizes 100, 200, and 300. Separate diffusion policy networks were trained for each dataset to investigate the effect of data quantity on performance. Each network underwent 300000 training steps. The evaluation was conducted on tasks with 4 or 6 objects. The initial UR3 end effector position was set above the bin. 202020 trials were conducted for each number of objects. In each trial, the robot’s performance was assessed until it either successfully completed the task (delivered all objects) or failed. Detailed data such as the number of grouping actions, grasping actions, the sequence of objects grasped per action, task completion status (i.e., whether all objects were successfully retrieved), failure reasons, and other related metrics were recorded. Table I shows the results of the trials and their statistics. The table contains numerous abbreviations. Their meanings are shown in the notes at the bottom. Especially in the statistical section, we computed the following values: (i) Completion rate: The rate at which all objects were successfully delivered over 20 trials. (ii) Delivery rate: The average number of successfully delivered objects divided by the initial number of objects across 20 trials. (iii) # Objects per grasp: The average number of objects delivered per successful grasp across 20 trials. These values conceptually reflect the Overall Success Rate (OSR) metric proposed by Chen et al. [29]. The maximum values of these statistics across all trained networks and numbers of objects are highlighted in lime, while the minimum values are highlighted in pink. When the value of column C is N, it indicates that the objects were not cleared. In this case, the detailed failure reason is listed in column R, which includes:(i) P: Pushing failure, where objects slipped from the gripper during pushing and failed to be grouped. (ii) G: Grasping failure, where the robot failed to grasp or dropped objects during delivery. (iii) Stagnation: The robot kept stationary or repeated the same action for more than 5 seconds. Stagnation during grouping was classified as a pushing failure, whereas stagnation during delivery was classified as a grasping failure. We can see from the results that as the number of randomly placed graspable objects increases, the task completion rates decrease. Failures frequently occur during pushing and grouping. The frequency of the failures increases with the number of initially randomly placed objects. The majority of pushing failures are caused by stagnation. This may be attributed to the increasing variety in the input images as the number of objects grows, making it more difficult for the neural network to determine the optimal pushing strategy based on limited information. We can also see from the results that the sequence of grouping and grasping is quite random. The network chose different policies according to object distributions. Furthermore, the results indicate that increasing the training data has a positive impact on all metrics. Notably, when the training data reaches 300300300, the robot grasps an average of 1.77 objects per attempt in the presence of 666 objects. The robot exhibits improved robustness during pushing and grouping and experienced fewer P failures. This demonstrates that the policy network can effectively support multi-object grasping and help improve picking efficiency.
Table II shows the detailed time costs when applying the policy trained on the dataset trained using 300 demonstrations to trials containing 6 objects. We can see from the results that while multi-object picking involves additional grouping operations that increase the average time per pick, the per-object efficiency improves significantly. In particular, the efficiency (time required per object) is reduced by approximately 8 seconds compared to single-object picking.
TABLE II: Time costs of 6 objects trained on the 300 dataset
ID $t_{1}$(s) $t_{2}$(s) $t_{3}$(s) T ID $t_{1}$(s) $t_{2}$(s) $t_{3}$(s) T
1 (12) (25) (26) 63 11 (11,28,13) - (26) 78
2 (32,13) (30,25) - 100 12 (30) - - -
3 (17,13) (18,18) - 66 13 - (12,35,21) - 68
4 - - - - 14 (12,55,13) - (22) 102
5 - (25) - - 15 (13) (27) (23) 63
6 (14,18,14) - (25) 71 16 - (18,20,22) - 60
7 (21) (24) (25) 70 17 - - - -
8 - (20,22,21) - 63 18 (15) (27) (23) 65
9 (13) (20) (20) 53 19 - - - -
10 (29) (25,22) - - 20 - (19,30,23) - 72
St Avg. time for 1 object:   19.3 s Efficiency for 1 object:   19.3 s
Avg. time for 2 objects:  22.9 s Efficiency for 2 objects:  11.4 s
Avg. time for 3 objects:  23.8 s Efficiency for 3 objects:   7.9 s
  • Note
    $t_{1}$ – Time required to pick and deliver a single object; $t_{3}$ – Time required to pick and deliver two objects simultaneously; $t_{3}$ – Time required to pick and deliver three objects simultaneously; $T$ – Total time required to clear all objects; Efficiency is computed by dividing the average time by the number of simultaneously picked objects.